\begin{thebibliography}{14}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi
\expandafter\ifx\csname bibnamefont\endcsname\relax
  \def\bibnamefont#1{#1}\fi
\expandafter\ifx\csname bibfnamefont\endcsname\relax
  \def\bibfnamefont#1{#1}\fi
\expandafter\ifx\csname citenamefont\endcsname\relax
  \def\citenamefont#1{#1}\fi
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\providecommand{\bibinfo}[2]{#2}
\providecommand{\eprint}[2][]{\url{#2}}

\bibitem[{\citenamefont{Einstein et~al.}(1935)\citenamefont{Einstein, Podolsky,
  and Rosen}}]{Einstein1935}
\bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Einstein}},
  \bibinfo{author}{\bibfnamefont{B.}~\bibnamefont{Podolsky}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{N.}~\bibnamefont{Rosen}},
  \bibinfo{journal}{Phys. Rev.} \textbf{\bibinfo{volume}{47}},
  \bibinfo{pages}{777} (\bibinfo{year}{1935}).

\bibitem[{\citenamefont{Bell}(1964)}]{Bell}
\bibinfo{author}{\bibfnamefont{J.~S.} \bibnamefont{Bell}},
  \bibinfo{journal}{Physics} \textbf{\bibinfo{volume}{1}}, \bibinfo{pages}{195}
  (\bibinfo{year}{1964}).

\bibitem[{\citenamefont{Hensen et~al.}(2015)\citenamefont{Hensen, Bernien,
  Dr\'eau, and et~al}}]{Hensen2015}
\bibinfo{author}{\bibfnamefont{B.}~\bibnamefont{Hensen}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Bernien}},
  \bibinfo{author}{\bibfnamefont{A.~E.} \bibnamefont{Dr\'eau}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{A.~R.} \bibnamefont{et~al}},
  \bibinfo{journal}{Nature} \textbf{\bibinfo{volume}{526}},
  \bibinfo{pages}{682} (\bibinfo{year}{2015}).

\bibitem[{\citenamefont{Genovese}(2005)}]{Genovese2005}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Genovese}},
  \bibinfo{journal}{Phys. Rep.} \textbf{\bibinfo{volume}{413}},
  \bibinfo{pages}{319} (\bibinfo{year}{2005}).

\bibitem[{\citenamefont{Sivia and Skilling}(2006)}]{Sivia2006}
\bibinfo{author}{\bibfnamefont{D.~S.} \bibnamefont{Sivia}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Skilling}},
  \emph{\bibinfo{title}{Data Analysis: A Bayesian Tutorial}}
  (\bibinfo{publisher}{Oxford University Press}, \bibinfo{year}{2006}).

\bibitem[{\citenamefont{Mathai and Provost}(1992)}]{Mathai1992}
\bibinfo{author}{\bibfnamefont{A.~M.} \bibnamefont{Mathai}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{S.~B.} \bibnamefont{Provost}},
  \emph{\bibinfo{title}{Quadratic forms in random variables}}
  (\bibinfo{publisher}{Taylor and Francis}, \bibinfo{year}{1992}).

\bibitem[{\citenamefont{Peres}(1978)}]{Peres1978}
\bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Peres}}, \bibinfo{journal}{Am.
  J. Phys.} \textbf{\bibinfo{volume}{46}}, \bibinfo{pages}{745}
  (\bibinfo{year}{1978}).

\bibitem[{\citenamefont{Lalo\"e}(2002)}]{Laloe2002}
\bibinfo{author}{\bibfnamefont{F.}~\bibnamefont{Lalo\"e}},
  \emph{\bibinfo{title}{Do we really understand Quantum Mechanics?}}
  (\bibinfo{publisher}{Cambridge University Press}, \bibinfo{year}{2002}).

\bibitem[{\citenamefont{Clauser et~al.}(1969)\citenamefont{Clauser, Horne,
  Shimony, and Holt}}]{Clauser1969}
\bibinfo{author}{\bibfnamefont{J.~F.} \bibnamefont{Clauser}},
  \bibinfo{author}{\bibfnamefont{M.~A.} \bibnamefont{Horne}},
  \bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Shimony}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{R.~A.} \bibnamefont{Holt}},
  \bibinfo{journal}{Phys. Rev. Lett.} \textbf{\bibinfo{volume}{23}},
  \bibinfo{pages}{880} (\bibinfo{year}{1969}).

\bibitem[{\citenamefont{Aerts}(1982)}]{Aerts1982}
\bibinfo{author}{\bibfnamefont{D.}~\bibnamefont{Aerts}},
  \bibinfo{journal}{Lettere al Nuovo Cimento} \textbf{\bibinfo{volume}{34}},
  \bibinfo{pages}{107} (\bibinfo{year}{1982}).

\bibitem[{\citenamefont{Warmuth and Kuzmin}(2009)}]{warmuth}
\bibinfo{author}{\bibfnamefont{M.~K.} \bibnamefont{Warmuth}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{D.}~\bibnamefont{Kuzmin}},
  \bibinfo{journal}{Machine Learning} \textbf{\bibinfo{volume}{78}},
  \bibinfo{pages}{63} (\bibinfo{year}{2009}).

\bibitem[{\citenamefont{Bush}(2015)}]{Bush2015}
\bibinfo{author}{\bibfnamefont{J.~W.~M.} \bibnamefont{Bush}},
  \bibinfo{journal}{Annu. Rev. Fluid Mech.} \textbf{\bibinfo{volume}{47}},
  \bibinfo{pages}{269} (\bibinfo{year}{2015}).

\bibitem[{\citenamefont{Varela et~al.}(1974)\citenamefont{Varela, Maturana, and
  Uribe}}]{VARELA}
\bibinfo{author}{\bibfnamefont{F.}~\bibnamefont{Varela}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Maturana}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Uribe}},
  \bibinfo{journal}{Biosystems} \textbf{\bibinfo{volume}{5}},
  \bibinfo{pages}{187} (\bibinfo{year}{1974}).

\bibitem[{\citenamefont{Trueblood and Busemeyer}(2011)}]{Trueblood2011}
\bibinfo{author}{\bibfnamefont{J.~S.} \bibnamefont{Trueblood}}
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{J.~R.}
  \bibnamefont{Busemeyer}}, \bibinfo{journal}{Cognitive Science}
  \textbf{\bibinfo{volume}{35}}, \bibinfo{pages}{1518} (\bibinfo{year}{2011}).

\end{thebibliography}

\documentclass[prl, twocolumn]{revtex4-1}
\usepackage{bm}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage[utf8]{inputenc}

\newcommand{\Tr}{\text{Tr}}

\begin{document}

\title{Fragile systems: A hidden variable theory for quantum mechanics}

\author{Yasmín Navarrete}
\affiliation{Université Grenoble Alpes}
\email{yasmin.navarrete@gipsa-lab.fr}

\author{Sergio Davis}
\affiliation{Comisión Chilena de Energía Nuclear}
\homepage{http://www.lpmd.cl/sdavis}
\email{sdavis@cchen.cl}

\date{\today}

\begin{abstract}
The formalism of Quantum Mechanics is derived from the application of Bayesian probability theory to ``fragile'' systems, i.e. systems that are perturbed by the act of measurement. Complex Hilbert spaces, non-commuting operators and the trace rule for expectations all arise naturally from the use of linear algebra to solve integral equations involving classical probabilities over hidden variables. We comment on the case of non-local hidden variables, where violations of Bell's theorem can be produced, as well as the non-fragile limit of the theory, where all measurements are commutative and the theory becomes analogous to classical statistical mechanics. 
\end{abstract}

\pacs{}

\keywords{}

\maketitle

%
%

\emph{Introduction}. \textemdash Quantum mechanics has been  a  controversial issue in  physics because of its non-local phenomena. 
In their seminal paper in 1935, Einstein, Podolsky and Rosen~\cite{Einstein1935} asserted that quantum mechanics was incomplete, because of the impossibility
  of predicting complementary quantities such as position and velocity of a particle at the same time. Moreover,  quantum entanglement introduced 
the notion of non-locality, which was later discussed by means of  Bell's theorem \cite{Bell}. Entangled systems manifest   non-classical 
correlations between outcomes performed on physical systems that are far apart,  but such that they have interacted in the past.

As a consequence of experimental violations of Bell's inequality~\cite{Hensen2015} it was concluded that quantum mechanics, if it is to be formulated in terms of hidden variables, has to be a non-local theory. In order to {attempt to fill this void in the understanding of the foundations of quantum mechanics, a number of hidden-variables theories~\cite{Genovese2005} have been  proposed.

In this {letter}, we will formulate a theory of fragile systems  based on non-local hidden variables derived from the application of Bayesian probability. We will recover the formalism of quantum theory from first principles, in particular, we will obtain that,

\begin{itemize}
    \item The  {states} after a measurement {correspond to fixed points of a linear transformation.}
    \item {This transformation leads to an eigenvalue equation involving a linear operator in Hilbert space}.
    \item The time evolution of {an arbitrary} state {is} described by a linear unitary operator.
    \item {Expectations are given by the trace rule of }the density matrix formalism.
\end{itemize}

%
%

\emph{Fragile systems}. \textemdash In simple terms, a fragile system is one that is affected by the act of observation (measurement). This distinguishes it from a non-fragile (classical) system, which is not modified upon observation. 

Because any system (being fragile or not) possesses information, we will think of a system as a ``black box'' 
that can be found in different \textbf{internal states}, to be denoted by $\lambda$. In general $\lambda$ 
contains many degrees of freedom, but we will not make use of that inner structure here. The internal state $\lambda$ 
contains all the information necessary to describe any aspect of the system.

We will consider a system with several real-valued, discrete observables $A$, $B$, $C$, \ldots. For instance, the observable 
$A$ may yield a value given by a real function $R_A(\lambda) \in \{a_1, \ldots, a_N\}$. In this case, the statement $a_k = R_A(\lambda)$ 
means that a measurement of $A$ when the system is found in its internal state $\lambda$ produced the value $a_k$.

The crucial difference between a fragile system and a non-fragile one is that, in a fragile system, \textbf{access 
to the internal state $\lambda$ is impossible}, because it is precisely this internal state which is modified 
by the act of measurement. We cannot, therefore, assume that we can evaluate $R_A$ on the internal state 
$\lambda$ to obtain the outcome of the measurement. As the modification of the state $\lambda$ depends on the 
details of the environment doing the measurement (which we do not know or control with accuracy), the outcome 
of a measurement is unavoidably stochastic, and a mathematical formulation requires probability theory. 

In summary, a fragile system is one that (a) modifies itself when it is measured, (b) when it is measured, the system remains in one state and  (c) its measurable properties have finite outcomes.

%
%

\emph{Probability theory}. \textemdash As we do not have exact knowledge of the internal state $\lambda$, we can only assign a probability distribution
over it, $P(\lambda|S)$, in \textbf{our state of knowledge} $S$. Unlike non-fragile systems, in a fragile system there is no state of 
knowledge $\mathcal{I}$ represented with an infinitely sharp peak, $P(\lambda|\mathcal{I})=\delta(\lambda-\lambda_0)$. Neither can we know the exact modification that a measurement will do on the internal state $\lambda$, thus for an observable $A$ we can only assign a transition probability 
$P(\lambda'|\lambda, A)$ of the final internal state $\lambda'$ given the initial state $\lambda$ and 
that a measurement of $A$ has occurred. 

By the application of the marginalization rule of probability~\cite{Sivia2006}, we see that if we are in a state of knowledge 
$S$ before a measurement of $A$ is made, after the measurement the new state of knowledge $S'$ will be given by

\begin{equation}
P(\lambda'|S') = \int d\lambda P(\lambda|S)P(\lambda'|\lambda, A)=P(\lambda'|S, A),
\label{eq_fragile_transform}
\end{equation}
so that $S'=S\wedge A$. In the particular case of a non-fragile system, the internal state $\lambda$ is not modified by the measurement of $A$, and therefore $P(\lambda'|\lambda, A)=\delta(\lambda'-\lambda)$, and we have thus $P(\lambda|S')=P(\lambda|S)$.

%
%

\emph{Fixed points of the transformation}. \textemdash Consider the situation after a measurement of $A$ yields the value $a_k$. Bayes' theorem~\cite{Sivia2006} tells us that our state of knowledge must agree with the probability

\begin{equation}
P(\lambda|a_k) = \frac{\delta(R_A(\lambda),a_k)P(\lambda|\mathcal{I}_0)}{P(a_k|\mathcal{I}_0)},
\label{eq_P_given_a}
\end{equation}
with
\begin{equation}
P(a_k|\mathcal{I}_0) = \int d\lambda P(\lambda|\mathcal{I}_0)\delta(R_A(\lambda),a_k).
\end{equation}

This implies our state is one of complete knowledge of $R_A$. We will postulate that the prior probability of the internal states $P(\lambda|\mathcal{I}_0)$ is flat, and then Eq. \ref{eq_P_given_a} reduces to

\begin{equation}
P(\lambda|a_k) = \frac{\delta(R_A(\lambda),a_k)}{\Omega_A(a_k)},
\label{eq_P_given_a_flat}
\end{equation}
where $\Omega_A(a_k)=\int d\lambda \delta(R_A(\lambda),a_k)$ is the density of internal states with given value of $R_A$. The fact that $P(\lambda|a_k)$ forbids all values of $\lambda$ with $R_A\neq a_k$ implies that two consecutive measurements of the same 
observable $A$, without any perturbation in between, will yield the same outcome $a_k$. From this it follows that the state of knowledge after a measurement must be a fixed point of the transformation $S \rightarrow S'$ given by Eq. \ref{eq_fragile_transform}. That is, if $g_k(\lambda)=P(\lambda|a_k)$ is the state of knowledge after obtaining the outcome $a_k$ in the measurement of $A$, we have

\begin{equation}
g_k(\lambda')=\int d\lambda g_k(\lambda)P(\lambda'|\lambda, A). 
\label{eq_fixed_point}
\end{equation}
\noindent

Obviously, this is always the case for a non-fragile system, as $P(\lambda'|\lambda,A)=\delta(\lambda'-\lambda)$.

%
%

\emph{Representation in terms of a complete basis}. \textemdash We can construct a complete, orthonormal basis $\{\phi_1(\lambda), \ldots, \phi_n(\lambda)\}$ for the probabilities $P(\lambda|S)$. Using the marginalization rule, 

\begin{equation}
P(\lambda|S) = \sum_{i=1}^N P(\lambda|a_k)P(a_k|S) = \sum_{i=1}^N \frac{\delta(R_A(\lambda),a_k)}{\Omega_A(a_k)}P(a_k|S).
\end{equation}

Now define the basis functions

\begin{equation}
\phi_i(\lambda) = \frac{\delta(R_A(\lambda),a_i)}{\sqrt{\Omega_A(a_i)}}
\end{equation}
such that

\begin{equation}
P(\lambda|S) = \sum_{i=1}^N v_i\phi_i(\lambda).
\label{eq_basis_expand_S}
\end{equation}

This fixes the coefficients $v_i=P(a_i|S)/\sqrt{\Omega_A(a_i)}$. Because the function $R_A(\lambda)$ is single-valued, $\phi_i(\lambda)\phi_j(\lambda)=0$ for any $\lambda$ if $i\neq j$. Furthermore, $\phi_i(\lambda)^2 = P(\lambda|a_i)$, so the basis is orthonormal,

\begin{equation}
\int d\lambda \phi_i(\lambda)\phi_j(\lambda) = \delta_{ij}.
\label{eq_ortho}
\end{equation}

Expanding also $P(\lambda'|S')$ in terms of this basis as

\begin{equation}
P(\lambda'|S') = \sum_{j=1}^n w_j\phi_j(\lambda'),
\label{eq_basis_expand_Sp}
\end{equation}
we can represent the states of knowledge $S$ and $S'$ by the vectors $\bm{v}=(v_1, \ldots, v_n)$ and 
$\bm{w}=(w_1, \ldots, w_n)$, respectively, and we can write Eq. \ref{eq_fragile_transform} as

\begin{equation}\label{eq_basis_expand2}
\sum_{j=1}^n w_j\phi_j(\lambda') = \int d\lambda \sum_{i=1}^n v_i\phi_i(\lambda)P(\lambda'|\lambda, A).
\end{equation}
\noindent
Multiplying both sides by $\phi_k(\lambda')$ and integrating over $\lambda'$, we have

\begin{eqnarray}
\sum_{j=1}^n w_j\int d\lambda' \phi_j(\lambda')\phi_k(\lambda') = \nonumber \\
\sum_{i=1}^n v_i\int d\lambda d\lambda' \phi_i(\lambda)P(\lambda'|\lambda, A)\phi_k(\lambda').
\end{eqnarray}

Now, using the orthonormality condition (Eq. \ref{eq_ortho}), 

\begin{equation}
\sum_{j=1}^n w_j\int d\lambda' \phi_j(\lambda')\phi_k(\lambda') = w_k,
\end{equation}
so we write 

\begin{equation}
w_k = \sum_{i=1}^n v_i\int d\lambda d\lambda' \phi_i(\lambda)P(\lambda'|\lambda, A)\phi_k(\lambda').
\end{equation}

Defining the matrix $\mathbb{T}$ with elements

\begin{equation}
T_{ij} = \int d\lambda d\lambda' \phi_i(\lambda')P(\lambda'|\lambda, A)\phi_j(\lambda).
\end{equation}
we can finally write Eq. \ref{eq_fragile_transform} as a linear transformation,

\begin{equation}
w_k = \sum_{i=1}^n T_{ki}v_i,
\end{equation}
equivalent to $\bm{w} = \mathbb{T}\cdot \bm{v}$. The fixed points of the transformation, namely $g_k(\lambda)$, are now 
encoded as eigenvectors $\bm{u}_k$ (with eigenvalue 1) such that $\bm{u}_k = \mathbb{T}\cdot \bm{u}_k$.

{The matrix} $\mathbb{T}$ is the transformation which allows us to obtain the fixed points of the system. On the other hand, we see that {we can obtain} an analogous operator $\mathbb{A}$ {leading to} the necessary transformation to obtain {also} the eigenvalues of the system {. For this} we instead use 

\begin{equation}
Q(\lambda') = \int d\lambda R_A(\lambda)P(\lambda|S)P(\lambda'|\lambda, A),
\end{equation}
such that for $S=a_k$, $P(\lambda|S)=P(\lambda|a_k)=g_k(\lambda)$. According to Eq. \ref{eq_P_given_a_flat}, $\lambda$ has 
zero probability if $R_A(\lambda)\neq a_k$. Considering this, we see that

\begin{equation}
Q(\lambda') = a_k\int d\lambda g_k(\lambda)P(\lambda'|\lambda, A) = a_k g_k(\lambda'),
\end{equation}
where the second equality holds because of Equation \ref{eq_fixed_point}. Finally,

\begin{equation}
a_k g_k(\lambda') = \int d\lambda R_A(\lambda) g_k(\lambda)P(\lambda'|\lambda,A)
\end{equation}
which, after performing the same basis expansion given by Eqs. \ref{eq_basis_expand_S} and \ref{eq_basis_expand_Sp}, becomes the eigenvalue problem 
\begin{equation}
a_i \bm{u}_i = \mathbb{A}\cdot \bm{u}_i.
\label{eq_eigenvalues}
\end{equation}

The matrix elements $A_{ij}$ are given by

\begin{equation}
A_{ij} = \int d\lambda d\lambda' R_A(\lambda)\phi_i(\lambda')P(\lambda'|\lambda, A)\phi_j(\lambda).
\label{eq_matrix_A}
\end{equation}

These matrix elements $A_{ij}$ are real numbers, because the function $R_A(\lambda)$ and the basis functions $\{\phi_i(\lambda)\}$ are real. However, in general it can be more convenient to express the eigenvalue problem in an arbitrary, complex basis $\{\psi_i(\lambda)\}$, so that 

\begin{equation}
\mathbb{A} = \sum_{i=1}^N a_i\bm{u}_i\otimes\bm{u}_i \rightarrow \sum_{i=1}^N a_i \bm{c}_i\otimes\bm{c}_i^*.
\end{equation}
with $\bm{c}_k$ a complex vector of dimension $N$, namely the coefficients of $\phi_k(\lambda)$ in the complex basis $\{\psi_i\}$. In this complex representation, the matrix $\mathbb{T}$ is unitary, and $\mathbb{A}$ is Hermitian.

%
%

\emph{Time evolution}. \textemdash Time evolution behaves also as a linear operator acting on states of knowledge,

\begin{equation}
P(\lambda'|S') = \int d\lambda P(\lambda|S) P(\lambda'|\lambda, \Delta t),
\end{equation}
with $S'=S\wedge \Delta t$. This operator is represented by the unitary matrix

\begin{equation}
U_{ij}(\Delta t) = \int d\lambda d\lambda' \phi_i(\lambda')P(\lambda'|\lambda, \Delta t)\phi_j(\lambda),
\end{equation}
such that an arbitrary state evolves as 

\begin{equation}
\bm{w}(t+\Delta t)=\mathbb{U}(\Delta t)\cdot \bm{w}(t).
\label{eq_time_evol_w}
\end{equation}
In quantum mechanics, this \emph{propagator} $\mathbb{U}(\Delta t)$ is given in terms of the Hamiltonian of the system as $\mathbb{U}(\Delta t)=\exp(-\mathrm{i}\mathbb{H}\Delta t/\hbar)$.

%
%

\emph{Density matrix formalism}. \textemdash In an arbitrary state of knowledge $S$ we can write the expectation value of the measurement $A$ as 

\begin{equation}
\Big<R_A\Big>_S = \int d\lambda R_A(\lambda)\sum_{i=1}^N P(\lambda|a_i)P(a_i|S)=\sum_{i=1}^N a_i P(a_i|S).
\end{equation}
\noindent
Now, recognizing that the $\{a_i\}$ are the eigenvalues of the matrix $\mathbb{A}$, as given by Eq. \ref{eq_eigenvalues}, with corresponding eigenvectors $\bm{u}_i$, we can write them as $a_i=(\bm{u}_i)^T\mathbb{A}\bm{u}_i$. {These are known as quadratic forms in Statistics~\cite{Mathai1992}}. Then, the expectation in state $S$ is

\begin{equation} \label{expect}
\Big<R_A\Big>_S = \sum_{i=1}^N p_i (\bm{u}_i)^T\mathbb{A}\bm{u}_i = \sum_{l,m} \rho_{ml} A_{lm} = \Tr \; (\mathbb{\rho A}),
\end{equation}
where the density matrix $\rho$ associated to the state of knowledge $S$ is defined as 

\begin{equation}
\rho = \sum_{i=1}^N p_i \bm{u}_i\otimes \bm{u}_i,
\end{equation}
with $p_i=P(a_i|S)$ and $\otimes$ the Kronecker product, $\bm{v}\otimes\bm{w}=\bm{v}\bm{w}^T$. This is a properly defined density matrix 
because the $p_i$ are probabilities of discrete propositions, non-negative and adding up to 1. {We can see that} every system {where} we can write the expectation value{s of its properties} in terms of the quadratic forms in Eq. (\ref{expect}) and {leading to} non-commutative {operators}, {can be considered as} a fragile system. 
It is straightforward to show, using Eq. \ref{eq_time_evol_w} on a particular basis expansion, that the density operator follows a von Neumann evolution, $\rho(t+\Delta t)=\mathbb{U}(\Delta t)\rho(t)\mathbb{U}^\dagger(\Delta t)$. 

%
%

\emph{Bell's Theorem} \textemdash In the present theory, the internal variables $\lambda$ play the role of a set of hidden variables for quantum mechanics. {Consider two systems $A$ and $B$ and} let $a$ and $a'$ detector settings on the system A, $b$ and $b'$ on the system B. The expectation values of the possible measurements for $A$ and $B$ with arbitrary settings $a$ and $b$ are given by

\begin{eqnarray}
\Big<R_A\Big>_{a,S}&=&\int d\lambda R_A(\lambda;a)P(\lambda|a, S),  \\ \nonumber
\Big<R_B\Big>_{b,S}&=&\int d\lambda R_B(\lambda;b)P(\lambda|b, S).
\end{eqnarray}

Taking the {joint} expected value for the system $A$ and $B$  for the {settings} $a$ and $b$ gives,

\begin{equation}
\Big<R_A R_B\Big>_{a,b,S} = \int d\lambda R_A(\lambda;a)R_B(\lambda;b)P(\lambda|a,b,S),
\end{equation}
which in general is different from $\big<R_A\big>_{a,S}\big<R_B\big>_{b,S}$ {because of the existence of correlations between the two systems}. Bell's inequality, {which in our notation reads}

\begin{widetext}
\begin{equation}
\Big| \Big<R_A R_B\Big>_{a,b,S}-\Big<R_A R_B\Big>_{a,a',S} \quad +\Big<R_A R_B\Big>_{b,b',S}+\Big<R_A R_B\Big>_{a',b',S}\Big| \leq 2.
\end{equation}
\end{widetext}
was applied in quantum mechanics in order to put some restrictions in its postulates on the possible hidden variables theories~\cite{Peres1978,Laloe2002,Clauser1969}. 

{In the case where} the probability distribution $P(\lambda|a,b,S)$ for the systems $A$ and $B$ {is not} separable {we have that} every  {observable} value depends on $\lambda$. {We have then}  a non-local hidden variable {theory} by construction.  Therefore the correlations of the systems $A$ and $B$ are non-local. 

{This shows that it is possible to have violations of Bell's inequality} in a macroscopic system  {if the corresponding} hidden variables are non-local, {and an interesting example of this behavior is given by Aerts\cite{Aerts1982}}. 

%
%

\emph{Conclusions}. \textemdash We have shown that fragile systems with discrete properties can be analyzed in terms of genuine quantum mechanics, complete with non-commuting operators and a density matrix formalism in complex Hilbert space. This not only gives a strong probabilistic justification for the fact that Nature itself is quantum mechanical, but it also opens the possibility of employing the structure of quantum mechanics as an inference tool in problems involving fragile systems outside physics, in areas such as biology, data analysis \cite{warmuth}, dynamical systems~\cite{Bush2015} among others. 
For instance, it would be possible  to apply this tool {to} biological models, under the perspective of autopoietic systems\cite{VARELA} having self-modifying properties, as we have shown that indeed  every system able to modify itself can be considered, under this formulation, as a fragile system. There is also the interesting possibility of applying our results as a formal justification of the recent idea of quantum cognition~\cite{Trueblood2011} in which the object of study is  human logic and human decisions. 

%
%

\emph{Acknowledgments}\textemdash SD gratefully acknowledges funding from FONDECYT 1140514, YN gratefully 
PhD fellowship of the Doctoral contract of the French government and Gipsa-lab.

\bibliography{fragile}
\bibliographystyle{apsrev}

\end{document}